OWG Meeting 26-Feb-2014

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm

Remote Connection

ESNet: 8542553

(if problems, call phone in conference room: 757-269-6460)

Previous Meeting


  • Announcements
  • Major systems reports
    • Trigger
    • DAQ
    • Monitoring
    • Controls


Attendees: Simon T., Yi Q., Mark I., Beni Z., Bryan M., Dmitry R., Mark D., David L.(chair), Sergey F., Graham H., Eugene C., Chris C., Sean D., Justin S., Hovanes E., Alex S., Sergey P.



  • Currently working on issue with FCAL DAQ where crate with the TI does not give signals
    • Alex has had some difficulty reproducing, but is working on it
  • Sergey is working on BCAL DAQ
    • First, is attempting from his personal account
    • Some issues using private CODA (or CODA in personal account?)
    • CODA 3.01 does not have the VME board drivers distributed with it
    • There looks to be one problematic board in the 2 BCAL crates tested so far
  • Chris C. reports that the GTP and SSP crates are in the hall but do not have CPUs yet.
    • Trunk lines are all in (expect for the Tagger)
    • Plan is to install TD <-> TI links first then fix the timing. This is as one-time calibration
    • BCAL will be done first since the boards have optical transceivers installed. FCAL will be next once optical transceivers are obtained and installed.
      • Waiting for 20 more optical transceivers, but may be able to borrow from TDs
    • Currently waiting for 1m cables to make connections (should be here any day)
  • Hovanes noted that the "Checkout Trigger" activity needs steps added in the project schedule


  • Different CODA versions will be used for different detector system-specific DAQs
    • 3.0 for FCAL
    • 3.01 for BCAL
    • 3.01 for the SoftROC testing facility
  • The Tagger fiber optics were installed going to the wrong place. There are a couple of options on how to fix this.


  • Work continues on the software ROC test facility that will be used for the 2014 Online Data Challenge


  • No meeting last week
  • Dave B.'s progress on interlocks is unknown
  • Hovanes is working on the EPICS support for FCAL
  • Orlando finished the BCAL pulser and can now start and stop it via CSS guy
  • The BCAL voltage controls are now live and will turn on/off voltages if the GUI is used
  • Nerses is coming around the 10th of March


Next Meeting

Next meeting is scheduled for March 12th